OpenStudy (anonymous):

y = cot2(sin θ) derivitive?

OpenStudy (anonymous):

and y = sin(tan 5x)

OpenStudy (anonymous):

Okay let \(u=\sin\theta\).\[ y=\cot^2(u) \]Now let \(v = \cot(u)\)\[ y=v^2 \]

OpenStudy (anonymous):

\[ \frac{dy}{d\theta }=\frac{dy}{dv}\frac{dv}{d\theta}=\frac{dy}{dv}\frac{dv}{du}\frac{du}{d\theta } \]

OpenStudy (anonymous):

\[\begin{split} \frac{dy}{d\theta} &=(2v)(-\csc^2(u))(\cos(\theta) )&v=\cot(u)\\ &=(2\cot (u))(-\csc^2(u))(\cos(\theta) )&u=\sin(\theta) \\ &=(2\cot (\sin\theta ))(-\csc^2(\sin \theta ))(\cos(\theta) ) \end{split} \]

OpenStudy (anonymous):

Chain rule can get a bit messy.